OpenStudy (anonymous):

9. x (years since 1991) 2 5 7 8 y (book sales billions $) 18 21 23 24 The table represents the book sales in the U.S in the years after 1991. Write an equation in slope-intercept form to represent the data. (Points : 1) y = x + 16 y = 3x + 18 y = 1/6 x + 16 y = 1/3 x + 18 10. What does the slope for this relation represent? x (years since 1991) 2 5 7 8 y (book sales billions $) 18 21 23 24 (Points : 1)

OpenStudy (blurbendy):

(21 - 18) = 3 (5 -2) = 3 y = x + 16 10) (21 - 18) = 3 (5 - 2) = 3 Therefore the slope is 1, and it represents the rate at which book sales have increased x years since 1991.